Could you please show me how to do matrix for these equations? I don't understand how it works.
7x+8y-z=9
x-2y-5z=-31
-7x+y+z=0

Thank you so very much for your time and help!
I really appreciate it!

Answer by Alan3354(21608)

Could you please show me how to do matrix for these equations? I don't understand how it works.
7x+8y-z=9
x-2y-5z=-31
-7x+y+z=0
-------------------
I can do that, but it's tedious to do on this site.
If you email me at gsihoutx@aol.com, I can scan and email you some hand-written notes, and an Excel sheet that I made that will solve 3x3 and 4x4 matrices.